OpenStudy (anonymous):

What is the equation of the line that passes through the points (1,2/3) and(-1,-1)?

OpenStudy (anonymous):

I will give you the basic steps. Tell me if you need me to help you further. First, find the slope between the two points. \[m=\frac{ y _{2}-y_{1} }{ x_{2}-x_{1} }\] Then, substitute the slope and one coordinate set of points into the slope-intercept form equation: \[y=mx+b\] Solve for b. Once you have m and b, you should be good to go. If you need it in standard form, I will assist you further if needed.

OpenStudy (anonymous):

I got the slope as 1 and 1/5 is that correct?

OpenStudy (anonymous):

One moment, I need to solve it myself.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

How did you get 1 1/5 for the slope?

OpenStudy (anonymous):

I plugged in the ys and the xs and then divided

OpenStudy (anonymous):

Good try, but unfortunately it is incorrect. Here is how to do it correctly: \[m=\frac{ -1-2/3 }{-1-1 }=\frac{ -1 and 2/3 }{-2 }\] -1 and 2/3 is simplified to -5/3 (mixed number to improper fraction) Now we are left with\[\frac{ -5/3 }{ -2 }\] We can divide them now. To divide fractions, we take the reciprocal of the second number, in this case -2 and make it 1/-2 (Whole numbers have a one as the denominator) then multiply. \[\frac{ -5 }{ 3 }*\frac{ 1 }{ -2 }\] After multiplying, we have -5/-6 or 5/6 m= 5/6 Now that we have this, can you do the second part?

OpenStudy (anonymous):

B= -1/6

OpenStudy (anonymous):

Correct. Good job. So, your equation is: y=5/6x-1/6

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

You're welcome.